OpenStudy (anonymous):

Solve, using the substitution method. m + n = 5 m – n = 3 A. There is no solution. B. The solution is (2, –1) C. The solution is (4, 1) D. There are an infinite number of solutions.

OpenStudy (anonymous):

i will give fan and medal o and a testominal

OpenStudy (campbell_st):

take the 2nd equation and make m the subject.. add n to both sides of the equation m = n + 3 now substitute this into the 1st equation for m (n + 3) + n = 5 or 2n + 3 = 5 solve for n when you get n, substitute the value into either equation and find m hope this helps